OpenStudy (anonymous):

4m + n = 26 m - 8m = 23 Solve by substution method I m lost s hek here

OpenStudy (anonymous):

is the second equation m-8m=23?

OpenStudy (anonymous):

what is the ? for I am lost here I will never get this right ugggg :/

OpenStudy (anonymous):

just asking if you wrote the question right... you wrote m-8m=23 isn't it m-8n=23 thats the question

OpenStudy (anonymous):

it is m - 8n = 23 thats how I see it

OpenStudy (anonymous):

oh ok i see my bd its late here in texas

OpenStudy (anonymous):

bad*

OpenStudy (anonymous):

substitution method means:write first one of the unknowns in terms of the other using either equations: 4m+n=26=>n=26-4m now substitute n=26-4m in the second equation: m-8(26-4m)=23=>m-208+32m=23=>33m=231=>m=7 now n=26-4m=26-4*7=26-28=-2